# Pre-trained Prompt-driven Community Search

### v1

@@ Abstract
The “pre-train, prompt” paradigm is widely adopted in various graph-based tasks and has shown promising performance in community detection. Most existing semi-supervised community detection algorithms detect communities based on known ones, and the detected communities typically do not contain the given query node. Therefore, they are not suitable for searching the community of a given node. Motivated by this, we adopt this paradigm into the semi-supervised community search for the first time and propose Pre-trained Prompt-driven Community Search (PPCS), a novel model designed to enhance search accuracy and efficiency. PPCS consists of three main components: node encoding, sample generation, and prompt-driven fine-tuning. Specifically, the node encoding component employs graph neural networks to learn local structural patterns of nodes in a graph, thereby obtaining representations for nodes and communities. Next, the sample generation component identifies an initial community for a given node and selects known communities that are structurally similar to the initial one as training samples. Finally, the prompt-driven fine-tuning component leverages these samples as prompts to guide the final community prediction. Experimental results on five real-world datasets demonstrate that PPCS performs better than baseline algorithms. It also achieves higher community search efficiency than semi-supervised community search baseline methods, with ablation studies verifying the effectiveness of each component of PPCS.
@@
Social networks, Semi-supervised community search, Graph pre-training, Graph prompt learning.
@@ I Introduction
Community structures commonly exist in networks, consisting of subsets of nodes that are closely connected [1, 2, 3]. Identifying these structures is crucial across various fields [4, 5, 6, 7] and gains significant attention in applications such as group recommendation [8, 9], fraud detection [10, 11], and event organization [12]. In real-world scenarios, prior information is often available, including pairwise constraints [13, 14, 15, 16, 17], node labels [18, 19], and known community structures [20, 21, 22, 23]. Semi-supervised community detection leverages such prior knowledge to enhance the identification of communities. Notably, some studies incorporate known communities as prior knowledge. For example, Bespoke [20] utilized community size and structure from training data to identify new communities, while SEAL [21] employs a generative adversarial network to detect and evaluate communities. Then, Wu et al. [22] proposed CLARE, which employs reinforcement learning to dynamically add or remove nodes, optimizing the accuracy of detected community structures. In addition, Wu et al. designed ProCom [23], a few-shot target community detection model that first captures latent community structures and leverages a few known samples to identify target communities.
@@
The above methods aim to detect communities similar to known ones. However, the detected communities typically do not contain the given query node, and some known communities exhibit structural discrepancies from the ground-truth community of a given node. As a result, the previously described methods are not applicable to identifying the given node’s community. To address this, semi-supervised community search focuses on identifying a given node’s community based on local information [24, 25, 26, 27]. For instance, Ni et al. [24] introduced SLSS, which utilizes structural similarity between communities to guide the identification of the given node’s community. Then, they further employed reinforcement learning to design SLRL [27], consisting of community extractors and expanders. Chen et al. [28] proposed CommunityAF, which integrates an autoregressive flow generation model with graph neural networks (GNN) to learn node embedding and generate communities. Although existing methods achieve promising results, they suffer from time-consuming community expansion due to subgraph similarity calculation in SLSS, redundant training in SLRL, and incremental GNN updates in CommunityAF. Unlike these approaches, our method reduces the number of community expansion steps and leverages a small number of training samples as prompts for efficient training and fast fine-tuning.
@@
Specifically, we apply the “pre-train, prompt” paradigm to the semi-supervised community search task and propose the Pre-trained Prompt-driven Community Search (PPCS) model. PPCS consists of three key components: node encoding, sample generation, and prompt-driven fine-tuning. Specifically, the node encoding component leverages GNN to learn the features of local structures in a graph, thereby obtaining the representation of nodes and communities. Then, for the sample generation component, an initial community is identified for a given node, and communities with similar structures are selected from known communities to form a specific training set for the given node. Finally, the prompt-driven fine-tuning component uses the specific training set as a prompt to predict the community of the given node. The main contributions of this paper are summarized as follows:
@@
    Motivated by the strong performance of the “pre-train, prompt” paradigm in community-related tasks, we introduce it to the semi-supervised community search task and propose PPCS. To the best of our knowledge, this is the first work to adopt the “pre-train, prompt” paradigm to the semi-supervised community search.
@@
    We propose PPCS, a “pre-train, prompt”-based community search method that reduces the number of community expansion steps and leverages a small number of training samples as prompts for efficient training and fast fine-tuning.
@@
    Extensive experiments on five real-world datasets demonstrate that PPCS outperforms baseline algorithms. It also achieves 37% higher community search efficiency compared to semi-supervised community search baseline methods, and ablation studies validate the effectiveness of each component in PPCS.
The rest of the paper is organized as follows. Section II summarizes the related works. Section III presents problem statement, PPCS overview, and each component’s detail of PPCS. Section IV reports experimental results. Section V concludes the paper.
@@ II Related works
This section will introduce related works on semi-supervised community detection, semi-supervised community search, and the “pre-train, prompt” paradigm for graph-based tasks.
@@ II-A Semi-supervised community detection
II-B Semi-supervised community search
Semi-supervised community search, known as semi-supervised local community detection, uses known community information to identify the community to which a given node belongs [24, 28, 27]. Ni et al. [24] first introduced the semi-supervised community search problem and proposed SLSS. SLSS leverages structural similarity between communities to guide the search for communities of given nodes. Ni et al. [27] further proposed SLRL, which incorporates reinforcement learning and consists of an extractor and an expander: the extractor captures the local structure of a given node and refines the training set via spectral clustering, while the expander derives the final community. Chen et al. [28] designed CommunityAF, which employs graph neural networks to learn node embeddings and uses an autoregressive flow generation model to detect communities containing given nodes. However, the above methods do not integrate the “pre-train, prompt” paradigm with the community search problem.
@@ II-B Semi-supervised community search
The “pre-train, prompt” paradigm was initially widely used in the field of natural language processing [30, 31]. In this paradigm, instead of modifying the pre-trained model for specific downstream tasks, this approach reformulates tasks through prompt design, aligning them with those addressed during pre-training. In recent years, the “pre-train and prompt” paradigm, known for its expressiveness and flexibility [32, 33], has also been widely adopted for graph-based tasks. For example, Sun et al. [34] proposed the All-in-one multi-task prompting method, which standardizes the format of the graph and language prompts, including prompt tokens, token structures, and insertion patterns, and then leverages prompts to adapt the pre-trained model to downstream tasks. Fang et al. [35] proposed GPF, a prompt-based general fine-tuning method for pre-training graph neural networks. Instead of manually designing a specific prompt function for each pre-training strategy, GPF adaptively generates prompt graphs for downstream tasks. Although there are works [23] on global community detection that model the task using the “pre-train, prompt” paradigm, this approach [23] is not suitable for semi-supervised community search.
@@ III Methodology
This section introduces the problem statement, the overview of the Pre-trained Prompt-driven Community Search (PPCS), and the details of the implementation of the PPCS components. Table I shows the notations used throughout this paper.
@@ III-A Problem statement
Given a graph $G=(V,E,A)$, a query node $v_{0}\in V$, and a set of known communities $C_{k}=\{c_{1},\cdots,c_{k}\}$, the goal of the semi-supervised community search problem is to predict the community $c_{v_{0}}$ of the query node $v_{0}$ using the communities in $C_{k}$.
III-B PPCS overview
Figure 1: Illustration of PPCS.
Following the “pre-train, prompt” paradigm, our model consists of three components: node encoding, sample generation, and prompt-driven fine-tuning. The detailed PPCS procedure is summarized in Algorithm 1 and illustrated in Figure 1. In the node encoding component (line 1), PPCS utilizes graph neural networks (GNN) and contrastive learning to generate node embeddings that capture the local topological features of nodes and their $k$-ego networks. These embeddings are subsequently leveraged to facilitate downstream community prediction. Once trained, the GNN encoder outputs node embeddings $z(v)\in R^{d}$ for each node $v\in V$.
@@ 4 Generate a final prediction community $c_{v_{0}}$;
Algorithm 1 PPCS
Finally, in the prompt-driven fine-tuning component (line 1), the model fine-tunes the prompt function $\text{PF}_{\phi}(\cdot)$ using the training samples, and then performs the final community prediction. First, PPCS initializes $\text{PF}_{\phi}(\cdot)$, implemented as a multilayer perceptron (MLP), using node embeddings obtained from the node encoding component. The prompt function is trained on the node embeddings of the sampled communities $C_{s}$, with its parameters $\phi$ optimized via a cross-entropy loss. The objective is to ensure that the nodes within the same community exhibit greater similarity in their local structural patterns with respect to their corresponding central nodes. After training, the $k$-ego network of the initial community, denoted as $N_{\dot{c}_{v_{0}}}$, is put into the prompt function to obtain the final prediction community: $c_{v_{0}}=\text{PF}_{\phi}(N_{\dot{c}_{v_{0}}})$ (line 1).
@@ III-C Node encoding
To model the relationship between nodes and their local structural context, PPCS aligns each node’s representation with that of its $k$-ego network. Specifically, the model randomly samples a batch of nodes $B\subset V$, and for each node $v\in B$, extracts its local structure $N_{v}$ consisting of $v$ and its $k$-hop neighbors. After GNN encoding, the model obtain the representations $z(v)$ and $z(N_{v})=\sum_{r\in N_{v}}z(r)$. The alignment is optimized using the loss from [23]:
@@
To enhance the encoder’s ability to capture community-relevant patterns, PPCS further discriminates and aligns local structural representations. For each node $v$, its $k$-ego local structure $N_{v}$ is perturbed by randomly removing nodes or edges, yielding a variant $\tilde{N}_{v}$. The encoder is then trained to align the embeddings of $N_{v}$ and $\tilde{N}_{v}$ using the following loss from [23]:
@@ 9      Update $\theta$ by applying gradient descent to minimize loss based on Equations (1) and (2);
PPCS finalize the node encoding by extracting both node-level and local structure-level features that reflect potential community structures. We input the graph $G=(V,E,A)$ into the trained encoder to obtain node embeddings $Z=\text{GNN}_{\theta}(E,A)$, where $z(v)\in Z$ denotes the initial embedding of node $v$. Community embeddings are computed via sum-pooling: $z(c)={\textstyle\sum_{v\in c}}z(v)$. These embeddings are then used for subsequent sample generation and prompt-driven fine-tuning.
@@ 12 Encode each training sample as $z(c^{i}),c^{i}\in C_{k}$;
In the sample generation component, the process consists of two steps. Step one: Based on reinforcement learning, PPCS uses known communities $C_{k}=\{c_{1},\cdots,c_{k}\}$ as the training set to train the agent. The trained agent then generates an initial community $\dot{c}_{v_{0}}$ for a given node $v_{0}$. Step two: our model computes the embedding similarity between the initial community and the known communities using Euclidean distance, and selects the top-$m$ most similar ones as samples.
@@ III-D1 Step one
During the reinforcement learning phase, PPCS adopts the policy gradient approach to optimize cumulative rewards. For each training community, a trajectory $\tau=(S_{0},a_{1},r_{1},\cdots,S_{t})$ is generated, where $S$, $a$, and $r$ represents the state, action, and reward. Specifically, each action is assigned a reward, which is defined as the change in F-score between $c_{t}$ and $c_{t-1}$ [22]:
@@ 7       Update $\phi$ by applying gradient descent to minimize loss based on Equation (10);
To enhance the efficiency of identifying the community of a query node, PPCS incorporates a prompt-driven fine-tuning mechanism [23]. This mechanism leverages a small number of similar communities as prompts to guide the extraction of the final community from the given node’s local structure.
Specifically, for a given query node $v_{0}$, PPCS first extracts a $k$-ego local structure $N_{\dot{c}_{v_{0}}}$ of initial community $\dot{c}_{v_{0}}$. The model then applies a prompt function $\text{PF}_{\phi}(\cdot)$ to identify the final prediction community $c_{v_{0}}$ from this local structure. The prompt function is implemented as a learnable multilayer perceptron (MLP) that estimates the similarity between $v_{0}$ and each node $u$ in $N_{\dot{c}_{v_{0}}}$. If this similarity exceeds a threshold $\alpha$, $u$ is added to $c_{v_{0}}$ [23]:
@@
To train the prompt function $\text{PF}_{\phi}(\cdot)$, PPCS utilizes a set of similar communities $C_{s}$ as supervision. For each community $c_{i}\in C_{s}$, a central node $w\in c_{i}$ is randomly selected, and its local structure $N_{w}$ is extracted. Each node $u\in N_{w},u\neq w$, is labeled based on whether it belongs to $c_{i}$. This binary label is used as the supervisory signal for training the MLP. The objective function is defined as [23]:
@@
In the community generation phase after model training, unlike previous semi-supervised community search methods [24, 27], PPCS does not rely on iterative community expansion. Instead, it directly identifies the final community using prompt-guided inference, thereby improving prediction efficiency.

### v2

# Pre-trained Prompt-driven Semi-supervised Local Community Detection

@@ Abstract
Semi-supervised local community detection aims to leverage known communities to detect the community containing a given node. Although existing semi-supervised local community detection studies yield promising results, they suffer from time-consuming issues, highlighting the need for more efficient algorithms. Therefore, we apply the “pre-train, prompt” paradigm to semi-supervised local community detection and propose the Pre-trained Prompt-driven Semi-supervised Local community detection method (PPSL). PPSL consists of three main components: node encoding, sample generation, and prompt-driven fine-tuning. Specifically, the node encoding component employs graph neural networks to learn the representations of nodes and communities. Based on representations of nodes and communities, the sample generation component selects known communities that are structurally similar to the local structure of the given node as training samples. Finally, the prompt-driven fine-tuning component leverages these training samples as prompts to guide the final community prediction. Experimental results on five real-world datasets demonstrate that PPSL outperforms baselines in both community quality and efficiency.
@@
Social networks, Semi-supervised local community detection, Graph pre-training, Graph prompt learning.
@@ I Introduction
Community structures are commonly observed in real networks [1, 2, 3]. Identifying community structures in networks has broad applications for various fields [4, 5, 6, 7], such as group recommendation [8, 9], fraud detection [10, 11], and event organization [12]. In real-world scenarios, network data often contains some prior information, such as pairwise constraints [13, 14, 15, 16, 17], node labels [18, 19], and known community structures [20, 21, 22, 23]. To improve the quality of community detection using such prior information, various semi-supervised community detection methods have been proposed [22]. For example, Bespoke [20] utilized the size and structure of known communities to identify new communities. SEAL [21] employs a generative adversarial network to extract a general community structure from known communities and generate new communities based on the extracted structure. Wu et al. [23] adopt “pre-train, prompt” paradigm to leverages a few known communities to identify communities.
@@
The above methods focus on detecting communities similar to the structure extracted from known communities. However, due to differences in community structures, the structure extracted from known communities may differ from the community structure of a given node. As a result, the detected communities may not include the given node. To address this, Ni et al. proposed a semi-supervised local community detection problem, which aims to identify the community of a given node using known communities [24, 25, 26, 27]. They further developed SLSS [24], leveraging structural similarity between communities to guide the identification of a given node’s community. Additionally, they introduced SLRL [27] based on reinforcement learning, which selects communities that are similar to the local structure of the given node to distill useful communities. Chen et al. [28] proposed CommunityAF, which integrates an autoregressive flow generation model with graph neural networks (GNN) to learn node embedding and generate communities. Although existing methods yield promising results, they suffer from time-consuming issues. These inefficiencies arise from frequent subgraph similarity calculations in SLSS, redundant training in SLRL, and frequent incremental GNN updates in CommunityAF. Consequently, there is a clear demand for more efficient algorithms.
@@
In response to this need, we apply the “pre-train, prompt” paradigm to the semi-supervised local community detection task and propose the Pre-trained Prompt-driven Semi-supervised Local community detection (PPSL) model. PPSL reduces the number of community expansion steps and leverages a small number of training samples as prompts for efficient training and fast fine-tuning. It consists of three key components: node encoding, sample generation, and prompt-driven fine-tuning. The node encoding component leverages GNN to learn the representation of nodes and communities, which is then used for both sample generation and prompt-driven fine-tuning. The sample generation component selects communities with structures similar to the local structure of the given node from known communities to form a training set. Finally, the prompt-driven fine-tuning component uses the training set as a prompt to predict the community of the given node.
The main contributions of this paper are summarized as follows:
@@
    Motivated by the strong performance of the “pre-train, prompt” paradigm in community-related tasks, we introduce it to the semi-supervised local community detection. To the best of our knowledge, this is the first work to adopt the “pre-train, prompt” paradigm to semi-supervised local community detection.
@@
    We propose a semi-supervised local community detection method based on “pre-train, prompt”, called PPSL. It reduces the number of community expansion steps and leverages a small number of training samples as prompts for efficient training and fast fine-tuning.
@@
    Extensive experiments on five real-world datasets demonstrate that PPSL outperforms the baseline algorithms in terms of community quality and also achieves faster runtime performance.
The rest of the paper is organized as follows. Section II summarizes the related works. Section III presents the problem statement, the PPSL overview, and each component of PPSL. Section IV reports experimental results. Section V concludes the paper.
@@ II Related works
This section will introduce related works on semi-supervised community detection, semi-supervised local community detection, and the “pre-train, prompt” paradigm for graph-based tasks.
@@ II-A Semi-supervised community detection
II-B Semi-supervised local community detection
Semi-supervised local community detection uses known community information to identify the community to which a given node belongs [24, 28, 27]. Ni et al. [24] first introduced the semi-supervised local community detection problem and proposed SLSS. SLSS leverages structural similarity between communities to guide the search for communities of given nodes. Ni et al. [27] further proposed SLRL, which incorporates reinforcement learning and consists of an extractor and an expander: the extractor captures the local structure of a given node and refines the training set via spectral clustering, while the expander derives the final community. Chen et al. [28] designed CommunityAF, which employs graph neural networks to learn node embeddings and uses an autoregressive flow generation model to detect communities containing given nodes. However, the above methods do not integrate the “pre-train, prompt” paradigm with the semi-supervised local community detection problem.
@@ II-B Semi-supervised local community detection
The “pre-train, prompt” paradigm was initially widely used in the field of natural language processing [30, 31]. In this paradigm, instead of modifying the pre-trained model for specific downstream tasks, this approach reformulates tasks through prompt design, aligning them with those addressed during pre-training. In recent years, the “pre-train and prompt” paradigm, known for its expressiveness and flexibility [32, 33], has also been widely adopted for graph-based tasks. For example, Sun et al. [34] proposed the All-in-one multi-task prompting method, which standardizes the format of the graph and language prompts, including prompt tokens, token structures, and insertion patterns, and then leverages prompts to adapt the pre-trained model to downstream tasks. Fang et al. [35] proposed GPF, a prompt-based general fine-tuning method for pre-training graph neural networks. Instead of manually designing a specific prompt function for each pre-training strategy, GPF adaptively generates prompt graphs for downstream tasks. Although there are works [23] on global community detection that model the task using the “pre-train, prompt” paradigm, this approach [23] is not suitable for semi-supervised local community detection.
@@ III Methodology
This section introduces the problem statement, the overview of the Pre-trained Prompt-driven Semi-supervised Local community detection (PPSL), and the details of the implementation of the PPSL components. Table I shows the notations used throughout this paper.
@@ III-A Problem statement
Given a graph $G=(V,E,A)$, a query node $v_{0}\in V$, and a set of known communities $C_{k}=\{c_{1},\cdots,c_{k}\}$, the goal of the semi-supervised local community detection problem is to predict the community $c_{v_{0}}$ of the query node $v_{0}$ using the communities in $C_{k}$.
III-B PPSL overview
Figure 1: Illustration of PPSL.
Following the “pre-train, prompt” paradigm, our model consists of three components: node encoding, sample generation, and prompt-driven fine-tuning. The detailed PPSL procedure is summarized in Algorithm 1 and illustrated in Figure 1. In the node encoding component (line 1), PPSL utilizes graph neural networks (GNN) and contrastive learning to generate node embeddings that capture the local topological features of nodes and their $k$-ego networks. These embeddings are subsequently leveraged to facilitate downstream community prediction. Once trained, the GNN encoder outputs node embeddings $z(v)\in R^{d}$ for each node $v\in V$.
@@ 4 Generate a final prediction community $c_{v_{0}}$;
Algorithm 1 PPSL
Finally, in the prompt-driven fine-tuning component (line 1), the model fine-tunes the prompt function $\text{PF}_{\phi}(\cdot)$ using the training samples, and then performs the final community prediction. First, PPSL initializes $\text{PF}_{\phi}(\cdot)$, implemented as a multilayer perceptron (MLP), using node embeddings obtained from the node encoding component. The prompt function is trained on the node embeddings of the sampled communities $C_{s}$, with its parameters $\phi$ optimized via a cross-entropy loss. The objective is to ensure that the nodes within the same community exhibit greater similarity in their local structural patterns with respect to their corresponding central nodes. After training, the $k$-ego network of the initial community, denoted as $N_{\dot{c}_{v_{0}}}$, is put into the prompt function to obtain the final prediction community: $c_{v_{0}}=\text{PF}_{\phi}(N_{\dot{c}_{v_{0}}})$ (line 1).
@@ III-C Node encoding
To model the relationship between nodes and their local structural context, PPSL aligns each node’s representation with that of its $k$-ego network. Specifically, the model randomly samples a batch of nodes $B\subset V$, and for each node $v\in B$, extracts its local structure $N_{v}$ consisting of $v$ and its $k$-hop neighbors. After GNN encoding, the model obtain the representations $z(v)$ and $z(N_{v})=\sum_{r\in N_{v}}z(r)$. The alignment is optimized using the loss from [23]:
@@
To enhance the encoder’s ability to capture community-relevant patterns, PPSL further discriminates and aligns local structural representations. For each node $v$, its $k$-ego local structure $N_{v}$ is perturbed by randomly removing nodes or edges, yielding a variant $\tilde{N}_{v}$. The encoder is then trained to align the embeddings of $N_{v}$ and $\tilde{N}_{v}$ using the following loss from [23]:
@@ 9      Update $\theta$ by applying gradient descent to minimize loss based on Equations (1) and (2);
PPSL finalize the node encoding by extracting both node-level and local structure-level features that reflect potential community structures. We input the graph $G=(V,E,A)$ into the trained encoder to obtain node embeddings $Z=\text{GNN}_{\theta}(E,A)$, where $z(v)\in Z$ denotes the initial embedding of node $v$. Community embeddings are computed via sum-pooling: $z(c)={\textstyle\sum_{v\in c}}z(v)$. These embeddings are then used for subsequent sample generation and prompt-driven fine-tuning.
@@ 12 Encode each training sample as $z(c^{i}),c^{i}\in C_{k}$;
In the sample generation component, the process consists of two steps. Step one: Based on reinforcement learning, PPSL uses known communities $C_{k}=\{c_{1},\cdots,c_{k}\}$ as the training set to train the agent. The trained agent then generates an initial community $\dot{c}_{v_{0}}$ for a given node $v_{0}$. Step two: our model computes the embedding similarity between the initial community and the known communities using Euclidean distance, and selects the top-$m$ most similar ones as samples.
@@ III-D1 Step one
During the reinforcement learning phase, PPSL adopts the policy gradient approach to optimize cumulative rewards. For each training community, a trajectory $\tau=(S_{0},a_{1},r_{1},\cdots,S_{t})$ is generated, where $S$, $a$, and $r$ represents the state, action, and reward. Specifically, each action is assigned a reward, which is defined as the change in F-score between $c_{t}$ and $c_{t-1}$ [22]:
@@ 7       Update $\phi$ by applying gradient descent to minimize loss based on Equation (10);
To enhance the efficiency of identifying the community of a query node, PPSL incorporates a prompt-driven fine-tuning mechanism [23]. This mechanism leverages a small number of similar communities as prompts to guide the extraction of the final community from the given node’s local structure.
Specifically, for a given query node $v_{0}$, PPSL first extracts a $k$-ego local structure $N_{\dot{c}_{v_{0}}}$ of initial community $\dot{c}_{v_{0}}$. The model then applies a prompt function $\text{PF}_{\phi}(\cdot)$ to identify the final prediction community $c_{v_{0}}$ from this local structure. The prompt function is implemented as a learnable multilayer perceptron (MLP) that estimates the similarity between $v_{0}$ and each node $u$ in $N_{\dot{c}_{v_{0}}}$. If this similarity exceeds a threshold $\alpha$, $u$ is added to $c_{v_{0}}$ [23]:
@@
To train the prompt function $\text{PF}_{\phi}(\cdot)$, PPSL utilizes a set of similar communities $C_{s}$ as supervision. For each community $c_{i}\in C_{s}$, a central node $w\in c_{i}$ is randomly selected, and its local structure $N_{w}$ is extracted. Each node $u\in N_{w},u\neq w$, is labeled based on whether it belongs to $c_{i}$. This binary label is used as the supervisory signal for training the MLP. The objective function is defined as [23]:
@@
In the community generation phase after model training, unlike previous semi-supervised local community detection methods [24, 27], PPSL does not rely on iterative community expansion. Instead, it directly identifies the final community using prompt-guided inference, thereby improving prediction efficiency.